Complex e. What is the ratio o to the married at each coll 2.2 SOC At St. Winnefred Hospital, the nun and females in various job categories are as follows. Totals 15 ince had the 178 163 Which society seer Nurse Orderly Lab technician Administrator 32 13 the percent change between 1997 and 2012 for ea Which states and pevisc had the largest increase and decrease? Which so ety seems to have the largest change in homicide 72 412 315 727 Totals r results in a paragraph Read each question carefully before constructing the fraction and solving for the answer. (HINT: Be State New Jersey lowa 8,053,000 2,852.000 4,139,000 19,439,000 32,268,000 3,074,186 26.059,203 426 1327 2579 1144 Texas California form Federal Bureau of Investigation al e Research: Approaches W. LAWRENCE NEUMA
media%2F640%2F640ab18b-686d-4a74-938f-3b
Answer #1

a) percentage of nurse which are male= 62/178*100 =    34.83%
  
b) proportion of female orderlies = 12/163 =   0.0736
  
c) for doctors,ratio of male to female = 83/15 or 83:15  
  
d)percentae of males =412/727*100=   56.67%
  
e) total male/total female = 412/315  
  
f) proportion of clerks that are male = 72/213=   0.3380
  
g) administrator = 13/727*100=   1.79
  
h) nurse to doctor = 178/98 = 89/49  
  
i) lab technician female to male = 30/32 = 15/16  
  
j) proportion of males that are clerk = 72/412=   0.1748